Find the number of degrees in each angle of a triangle if the first angle is twice the second and three times the third in value.
Answer by midwood_trail(310)

Find the number of degrees in each angle of a triangle if the first angle is twice the second and three times the third in value.
====================================================
The total measure of the angles in any triangle = 180 degrees.
Each triangle has 3 angles.
Here are the 3 angles:
first angle = 2x
second angle = x
third angle = 3x
2x + x + 3x = 180
6x = 180
x = 180/6
x = 30
Now, replace x with 30 for each given angle.
first angle = 2x = 2(30) = 60
=============================
second angle = x = 30
=============================
third angle = 3x = 3(30) = 90
Did you follow?